# MMAD: Multi-label Micro-Action Detection in Videos

### v2

@@ 1 Introduction
Human body actions, as an important form of non-verbal communication, effectively convey emotional information in social interactions [2]. Previous research primarily focused on interpreting classical expressive emotions through facial expressions [17, 54, 39, 10], speech [12, 60, 19, 32], or expressive body gestures [3, 35, 8, 23, 22]. In contrast, our study shifts the focus to a specific subset of body actions known as Micro-Actions (MAs) [35, 8, 14, 15, 13, 24]. MAs are imperceptible non-verbal behaviors characterized by low-intensity movement with potential applications in emotion analysis.
@@
Compared to traditional actions [46, 21, 7, 26], MAs have distinct characteristics as follows: (1) Short duration. As shown in Fig. 1, MAs typically last only a few seconds, exhibiting subtle visual changes between consecutive frames. For instance, “touching neck” only exhibits minor changes in the neck region between a few frames. In contrast, conventional actions typically last around 5–10 seconds and involve larger and more dynamic motions in hundreds of frames. For example, the movements of “archery” or “jump” involve large motions. (2) Low intensity. MAs are characterized by minor spatial distinctions. As shown in Fig. 1 (b), the difference between “touching neck” and “touching shoulder” varies only in the specific contact regions involved. In contrast, conventional actions usually with identifiable motion patterns, such as those in “longjump” in Fig. 1 (a), where the overall movement is more visually distinct. (3) Fine-grained categories. MAs demand classification at both the body-part and action levels, involving isolated movements of individual body parts (e.g., “head”, “upper limb”, and “lower limb”) as well as coordinated motions combining parts (e.g., “head-hand,” and “body-hand”). In contrast, conventional action recognition typically focuses on larger-scale, whole-body motions.
Remarkable progress has been made in the micro-action recognition [14, 24] task with the advance of Vision Transformers [13, 52, 25]. However, in the real world, micro-actions are naturally temporal co-occurring, which poses challenges for traditional action recognition methods. As shown in Fig. 1 (c), different micro-actions may occur simultaneously, such as “stretching arms” frequently happening simultaneously with “putting hands together”. Therefore, driven by this intuition, we propose a new task named Multi-label Micro-Action Detection (MMAD) that recognizes all the micro-actions in the video sequence, achieving a fine-grained understanding of micro-actions. MMAD involves identifying all micro-actions within the video and determining their corresponding start and end times, as well as their categories. Firstly, MMAD requires a model capable of capturing both long-term and short-term action relationships to locate multi-scale micro-actions. Secondly, the model also needs to explore the complex inter-relationships between different micro-actions to ensure comprehensive detection of all possible micro-actions. Finally, due to the inherent nature of short duration and subtle movements in micro-actions, there is also a greater challenge in recognizing the correct categories.
@@
  • We evaluate 10 baselines from the conventional temporal action detection on MMAD and in-depth studies, which reveals the inherent challenges in multi-label micro-action detection.
@@ 2.1 Micro-Actions Recognition
Micro-Actions (MAs) [35, 8, 14] are an important form of non-verbal communication, which is usually related to humans’ emotional status [2]. To facilitate the study of these subtle movements, several datasets have been constructed. To advance the study of these subtle movements, several datasets have been developed. iMiGUE [35] and SMG [8] focused on spontaneous micro-gestures in the upper limbs of athletes, revealing deep emotional states conveyed through these micro-gestures. In contrast, MPIIGI [3] primarily examined subtle upper-body behaviors in group interactions. To better analyze and understand the whole-body movement, Guo et al. [14] proposed a large-scale micro-action dataset named Micro-Action (MA-52), which consists of 52 action-level MAs within 7 body-level in whole-body. They also evaluated conventional action recognition methods, including 2D CNN based [53, 29, 11], 3D CNN based [50, 6], GCN-based [58, 40], and Transformer-based [41]. More recently, Li et al. [24] proposed a prototypical calibrating ambiguous network, designed to mitigate the influence of the inherent ambiguity of micro-actions in micro-action recognition.
@@ 3.1 Dataset Construction
Data Collection. The proposed Multi-label Micro-Action-52 (MMA-52) dataset is built upon the MA-52-Pro dataset collected by [14]. However, MA-52-Pro is not directly applicable to micro-action detection tasks due to the following reasons: 1) Lack of fine-grained annotations: MA-52-Pro does not provide detailed start/end timestamps for individual action instances. 2) Variation in video lengths: The videos in MA-52-Pro vary significantly in length, each video contains 1 to 15 MAs with duration ranging from 5s to beyond 100s. Such substantial imbalance in the action instances makes it unsuitable for action detection tasks. To address these challenges, first, we segment the videos into smaller clips, each ranging from 5 to 15 seconds, to ensure more consistency in video sequence and action instances. Next, we annotate each micro-action instance with its corresponding categories and precise start/end timestamps.
Figure 3: The ratio of each action instance category in the MMA-52 dataset. “A1, A2, …, G4” denote the action-level categories while “A, B, …, G” denote the body-level categories. The distribution of micro-action instances across the action-level and body-level follows a long-tail pattern.
@@ 3.2 Dataset Statistics and Properties
Based on the above data statistics, the characteristics of the proposed MMA-52 dataset can be summarized as follows. 1) Long-tail category distribution. As shown in Fig. 3, the dataset exhibits a long-tail distribution, where some MAs occur frequently while others rarely occur. For example, “B3: Turning head” accounts for nearly 13% whereas “A2: Turning around” appears not up to 1%. This imbalance presents a challenge for models to enhance their robustness and ability to generalize in rare MAs. 2) Variability in micro-action hierarchies. As shown in Fig. 4, micro-action duration exhibits significant variability, particularly at the action level, where some actions (e.g., C3, D4, E6) last considerably longer than others. The high standard deviation suggests substantial fluctuations in duration across individuals and scenarios, posing challenges for model learning and prediction. In contrast, the body-level shows a relatively lower standard deviation, indicating that body-level labels are more stable. This highlights the need for models to capture hierarchical relationships, as individual actions are influenced by different body parts, making it insufficient to rely solely on global features. 3) Subject-independent evaluation. Given the subtle nature of micro-actions, and micro-action patterns will be different across individuals, the MMA-52 dataset adopts a subject-independent setting. This means that the dataset includes action instances from a diverse range of subjects, ensuring that there is no overlap of subjects between the training, validation, and test sets. The goal is to ensure that the model learns to identify and generalize micro-actions across varying body types and movements.
@@ 3.3 Comparison with Existing Datasets
First, we review the related datasets in micro-action recognition, e.g., PAVIS F-T [4], iMiGUE [35], and SMG [8]. PAVIS F-T [4] and MPIIGI [3] focus on body behaviors in group social interactions. The former concentrates solely on face-touching versus non-face-touching, while the latter analyzes a broader range of behavior categories (e.g., scratch and shrug). In contrast, iMiGUE [35] and SMG [8] target upper limb micro-gestures. However, these datasets are relatively limited in terms of category diversity and subjects. To address this limitation, MA-52 [14] introduces the large-scale micro-action recognition dataset with 52 categories. Then, we compare with micro-action detection datasets. iMiGUE and SMG can be applied to action detection tasks, but they are unavailable due to privacy concerns. In contrast, our MMA-52 benefits from the hierarchy label (i.e., body-level and action-level), large-scale action instances involving diverse subjects, enabling the design of more comprehensive and scalable detection models.
@@ 4.3 Dual-path Spatial-Temporal Adapter
where [;][;][ ; ] denotes the concatenate operation, $\beta$ and $\gamma$ are two learnable parameters to balance the weights of spatial and temporal pathways, respectively. S-DWConv symbols the spatial depth-wise convolution with the kernel size of $1\times 1$.
@@ 5.1 Experiments Setup
Evaluation Metrics. We use the Detection-mAP [47] to evaluate the performance of multi-label micro-action detection. Detection-mAP measures the completeness of predicted action instances.
We report the average mAP results for tIoU thresholds ranging from 0.1 to 0.9 in increments of 0.1 and the average mAP at each specific threshold. Considering the hierarchy of micro-actions, we report Detection-mAP at both body-level and action-level. Additionally, we also report the average value of Detection-mAP (AVG) of these two levels.
Implementation Details. We conduct experiments with the open-source toolbox OpenTAD [34]. The model employs mixed-precision training and activation checkpointing to reduce memory usage. Following [33], we use ActionFormer [61] as the detector head, retaining the original hyperparameter settings. The backbone’s learning rate remains fixed, while the adapter’s learning rate varies between 1e-4 and 4e-4. By default, frame resolution is set to 1602. More implementation details are attached in the Appendix.
@@ 5.2 Main Results
Since Multi-label Micro-Action Detection (MMAD) is a new task, we evaluate 10 baselines in conventional temporal action detection. The results are reported in Table 3. (1) For the multi-label temporal action detection methods (MS-TCT [9] and PointTAD [47]), these methods achieve the lowest performance, we attribute it that there is a gap between conventional actions and micro-actions, which leads to performance significantly drop. (2) For the feature-based methods, TemporalMaxer [48] get the best average mAP of 20.34, while the TriDet [45] only achieves the average mAP of 16.04. (3) For the feature-based methods, the baseline method AdaTAD [34] exhibits better performance with the backbone model enlarged from the VideoMAE small version to the base version. The best result in average mAP is 21.08, which is better than all feature-based methods. Compared to the baseline, we can see that the proposed method exhibits consistent improvement on different backbones. Specifically, on the VideoMAE-S and VideoMAE-B, there are 1.63%, and 2.26% improvements on average mAP, respectively. Although the proposed method achieves the highest average mAP of 25.34, it is still far away from conventional action detection [33]. These results indicate that there is still a gap in accurately identifying the micro-actions.
@@
  Setting Param. D-mAP gains
@@ 5.3 Ablation Studies
The ablation of the adapters. To validate the effectiveness of the proposed dual-path spatio-temporal adapter, we conduct experiments as follows: “Full fine-tuning” denotes fine-tuning the backbone, Standard Adapter [18], and “TIA” from the baseline model AdaTAD. The results are reported in Table 4. The baseline model only achieves the D-mAP of 16.25%, due to the neglect of curial spatial information in micro-actions. In contrast, the proposed DSTA achieves the best results of 18.16 in terms of D-mAP, and there are 1.91% improvements.
The ablation of parameters $\beta$ and $\gamma$. As shown in Figure 5, we conduct experiments to evaluate the impact of learnable parameters $\beta$ and $\gamma$ in spatial-path and temporal-path, respectively. As shown in Table 5, there is a performance drop in both “w/o $\beta$” and “w/o $\gamma$”. For example, 18.16 to 15.90 in “w/o $\beta$”, and 18. 16 to 16.08 in “w/o $\gamma$” in terms of mAP. The results show that the performance drop in “w/o $\beta$” is larger than “w/o $\gamma$”, highlighting the importance of spatial information in identifying the micro-actions.
Table 5: Ablation on parameter $\beta$ and $\gamma$.
  Action-level Detection-mAP
Models @0.2 @0.5 @0.7 Avg.
w/o $\beta$ 24.82 17.07 9.05 15.90
w/o $\gamma$ 29.20 17.20 9.05 16.08
w/o $\beta$ w/o $\gamma$ 27.09 19.64 10.31 17.30
Full Model 28.05 20.40 9.02 18.16

### v3

@@ 1 Introduction
Human body actions, as an important form of non-verbal communication, effectively convey emotional information in social interactions [2]. Previous research primarily focused on interpreting classical expressive emotions through facial expressions [20, 66, 47, 11], speech [23, 39, 77], or expressive body gestures [3, 43, 9, 27, 26]. In contrast, our study shifts the focus to a specific subset of body actions known as Micro-Actions (MAs) [43, 9, 40, 17, 18, 15, 29]. MAs are imperceptible non-verbal behaviors characterized by low-intensity movement with potential applications in emotion analysis.
@@
Compared to traditional actions [55, 25, 7, 12, 33, 30, 64, 13], MAs have distinct characteristics as follows: (1) Short duration. As shown in Fig. 1, MAs typically last only a few seconds, exhibiting subtle visual changes between consecutive frames. For instance, “touching neck” only exhibits minor changes in the neck region between a few frames. In contrast, conventional actions typically last around 5–10 seconds and involve larger and more dynamic motions in hundreds of frames. For example, the movements of “archery” or “jump” involve large motions. (2) Low intensity. MAs are characterized by minor spatial distinctions. As shown in Fig. 1 (b), the difference between “touching neck” and “touching shoulder” varies only in the specific contact regions involved. In contrast, conventional actions usually with identifiable motion patterns, such as those in “longjump” in Fig. 1 (a), where the overall movement is more visually distinct. (3) Fine-grained categories. MAs demand classification at both the body-part and action levels, involving isolated movements of individual body parts (e.g., “head”, “upper limb”, and “lower limb”) as well as coordinated motions combining parts (e.g., “head-hand,” and “body-hand”). In contrast, conventional action recognition typically focuses on larger-scale, whole-body motions.
Remarkable progress has been made in the micro-action recognition [17, 29, 15, 61, 32] task with the advancement of Vision Transformers [28, 63, 62, 68, 31, 67]. However, in the real world, micro-actions are naturally temporal co-occurring, which poses challenges for traditional action recognition methods. As shown in Fig. 1 (c), different micro-actions may occur simultaneously, such as “stretching arms” frequently happening simultaneously with “putting hands together”. Therefore, driven by this intuition, we propose a new task named Multi-label Micro-Action Detection (MMAD) that recognizes all the micro-actions in the video sequence, achieving a fine-grained understanding of micro-actions. MMAD involves identifying all micro-actions within the video and determining their corresponding start and end times, as well as their categories. Firstly, MMAD requires a model capable of capturing both long-term and short-term action relationships to locate multi-scale micro-actions. Secondly, the model also needs to explore the complex inter-relationships between different micro-actions to ensure comprehensive detection of all possible micro-actions. Finally, due to the inherent nature of short duration and subtle movements in micro-actions, there is also a greater challenge in recognizing the correct categories.
@@
  • We evaluate 10 baselines from the conventional temporal action detection on MMAD and in-depth studies, which reveal the inherent challenges in multi-label micro-action detection.
@@ 2.1 Micro-Actions Recognition
Micro-Actions (MAs) [43, 9, 17, 8, 29] are an important form of non-verbal communication, which are usually related to humans’ emotional status [2]. To facilitate the study of these subtle movements, several datasets have been constructed. To advance the study of these subtle movements, several datasets have been developed. iMiGUE [43] and SMG [9] focused on spontaneous micro-gestures in the upper limbs of athletes, revealing deep emotional states conveyed through these micro-gestures. In contrast, MPIIGI [3] primarily examined subtle upper-body behaviors in group interactions. To better analyze and understand the whole-body movement, Guo et al[17] proposed a large-scale micro-action dataset named Micro-Action (MA-52), which consists of 52 action-level MAs within 7 body-level in whole-body. They also evaluated conventional action recognition methods, including 2D CNN based [65, 36, 14], 3D CNN based [59, 6], GCN-based [72, 48], and Transformer-based [49]. More recently, Li et al[29] proposed a prototypical calibrating ambiguous network, designed to mitigate the influence of the inherent ambiguity of micro-actions in micro-action recognition.
@@ 3.1 Dataset Construction
Data Collection. The proposed Multi-label Micro-Action-52 (MMA-52) dataset is built upon the MA-52-Pro dataset collected by [17]. However, MA-52-Pro is not directly applicable to micro-action detection tasks due to the following reasons: 1) Lack of fine-grained annotations: MA-52-Pro does not provide detailed start/end timestamps for individual action instances. 2) Variation in video lengths: The videos in MA-52-Pro vary significantly in length, each video contains 1 to 15 MAs with durations ranging from 5s to beyond 100s. Such substantial imbalance in the action instances makes it unsuitable for action detection tasks. To address these challenges, first, we segment the videos into smaller clips, each ranging from 5 to 15 seconds, to ensure more consistency in video sequence and action instances. Next, we annotate each micro-action instance with its corresponding categories and precise start/end timestamps.
Figure 3: The ratio of each action instance category in the MMA-52 dataset. “A1, A2, …, G4” denote the action-level categories while “A, B, …, G” denote the body-level categories. The distribution of micro-action instances across the action-level and body-level follows a long-tail pattern. The definition of each category is the same as that of the MA-52 dataset [17].
@@ 3.2 Dataset Statistics and Properties
Based on the above data statistics, the characteristics of the proposed MMA-52 dataset can be summarized as follows. 1) Long-tail category distribution. As shown in Fig. 3, the dataset exhibits a long-tail distribution, where some MA occur frequently while others rarely occur. For example, “B3: Turning head” accounts for nearly 13% whereas “A2: Turning around” appears not up to 1%. This imbalance presents a challenge for models to enhance their robustness and ability to generalize in rare MAs. 2) Variability in micro-action hierarchies. As shown in Fig. 4, micro-action duration exhibits significant variability, particularly at the action level, where some actions (e.g., C3, D4, E6) last considerably longer than others. The high standard deviation suggests substantial fluctuations in duration across individuals and scenarios, posing challenges for model learning and prediction. In contrast, the body-level shows a relatively lower standard deviation, indicating that body-level labels are more stable. This highlights the need for models to capture hierarchical relationships, as individual actions are influenced by different body parts, making it insufficient to rely solely on global features. 3) Subject-independent evaluation. Given the subtle nature of micro-actions, and micro-action patterns will be different across individuals, the MMA-52 dataset adopts a subject-independent setting. This means that the dataset includes action instances from a diverse range of subjects, ensuring that there is no overlap of subjects between the training, validation, and test sets. The goal is to ensure that the model learns to identify and generalize micro-actions across varying body types and movements.
@@ 3.3 Comparison with Existing Datasets
We first review the related datasets [4, 3, 43, 9] in micro-action recognition. PAVIS F-T [4] and MPIIGI [3] focus on body behaviors in group social interactions. The former concentrates solely on face-touching versus non-face-touching, while the latter analyzes a broader range of behavior categories (e.g., scratch and shrug). In contrast, iMiGUE [43] and SMG [9] target upper limb micro-gestures. However, these datasets are relatively limited in terms of category diversity and subjects. To address this limitation, MA-52 [17] introduces a large-scale micro-action recognition dataset with 52 categories. Then, we compare with micro-action detection datasets. iMiGUE and SMG can be applied to action detection tasks, but they are limited to action categories. In contrast, our MMA-52 benefits from the hierarchy label (i.e., body-level and action-level), large-scale action instances involving diverse subjects, enabling the design of more comprehensive and scalable detection models.
@@ 4.3 Dual-path Spatial-Temporal Adapter
where $[;]$ denotes the concatenation operation, $\beta$ and $\gamma$ are two learnable parameters to balance the weights of spatial and temporal pathways, respectively. S-DWConv symbols the spatial depth-wise convolution with the kernel size of $1\times 1$.
@@ 5.1 Experiments Setup
Evaluation Metrics. We use the mean Average Precision (mAP) [56, 74, 41] to evaluate the performance of multi-label micro-action detection. mAP measures the completeness of predicted action instances. We report the average mAP results for tIoU thresholds ranging from 0.1 to 0.9 in increments of 0.1 and the average mAP at each specific threshold. Considering the hierarchy of micro-actions, we report Detection-mAP at both body-level and action-level. Additionally, we also report the average value of mAP (AVG) of these two levels.
Implementation Details. We conduct experiments with the open-source toolbox OpenTAD [42]. The model employs mixed-precision training and activation checkpointing to reduce memory usage. Following [41], we use ActionFormer [74] as the detector head, retaining the original hyperparameter settings. The backbone’s learning rate remains fixed, while the adapter’s learning rate varies between 1e-4 and 4e-4. By default, frame resolution is set to 1602. For the body-level prediction, we follow the common practice in micro-action analysis [17, 16] by converting the action-level results to the body-level.
@@ 5.2 Main Results
Since Multi-label Micro-Action Detection (MMAD) is a new task, we evaluate 10 baselines in conventional temporal action detection. The results are reported in Table 3. (1) For the multi-label temporal action detection methods (MS-TCT [10] and PointTAD [56]), these methods achieve the lowest performance. We attribute the significant performance drop to the gap between conventional actions and micro-actions. (2) For the feature-based methods, TemporalMaxer [57] get the best average mAP of 20.34, while the TriDet [54] only achieves the average mAP of 16.04. (3) For the feature-based methods, the baseline method AdaTAD [42] exhibits better performance with the backbone model enlarged from the VideoMAE small version to the base version. The best result in average mAP is 21.80, which is better than all feature-based methods. Compared to the baseline, we can see that the proposed method exhibits consistent improvement on different backbones. Specifically, on the VideoMAE-S and VideoMAE-B, there are 1.63%, and 2.26% improvements on average mAP, respectively. Although the proposed method achieves the highest average mAP of 25.34, it is still far away from conventional action detection [41]. These results indicate that there is still a gap in accurately identifying the micro-actions.
@@
  Setting Param. mAP gains
@@ 5.3 Ablation Studies
The ablation of the adapters. To validate the effectiveness of the proposed dual-path spatio-temporal adapter, we conduct experiments as follows: “Full fine-tuning” denotes fine-tuning the backbone, Standard Adapter [21], and “TIA” from the baseline model AdaTAD. The results are reported in Table 4. The baseline model only achieves the mAP of 16.25%, due to the neglect of crucial spatial information in micro-actions. In contrast, the proposed DSTA achieves the best results of 18.16 in terms of mAP, and there is 1.91% improvement.
@@ 6 Conclusions
Acknowledgments
This work is supported by National Key R&D Program of China (NO.2024YFB3311602), Natural Science Foundation of China (62272144), the Anhui Provincial Natural Science Foundation (2408085J040), Anhui Provincial Key Research and Development Project (202304a05020068), the Major Project of Anhui Provincial Science and Technology Breakthrough Program (202423k09020001), the Fundamental Research Funds for the Central Universities (JZ2024HGTG0309, JZ2024AHST0337), and the Earth System Big Data Platform of the School of Earth Sciences, Zhejiang University.
Appendix A Dataset Details
The Multi-label Micro-Action-52 (MMA-52) dataset is designed for multi-label micro-action detection. Each action instance contains both Body-level and Action-level categories. Body-level category denotes the body part of micro-action occurring, including [ “A: Body”, “B: Head”, “C: Upper limb”, “D: Lower limb”, “E: Body-hand”, “F: Head-hand”, and “G: Leg-hand”.] Action-level category denotes the exact name of micro-actions. Taking the body-level category “A: Body” as an example, there are 5 action-level categories: “A1: Shaking body”, “A2: Turning around”, “A3: Sitting straightly”, “A4: Shrugging” and “A5: Rising up”. Body-level labels and action-level labels are naturally hierarchical structures. In summary, there are 7 Body-level categories and 52 Action-level categories. The detailed descriptions and label ID of each micro-action are the same as the MA-52 dataset [17].
Appendix B Implementation details
Figure 9: Overview of the proposed Dual-path Spatial-Temporal Adapter.S-DWConv” denotes the spatial-path while “T-DWConv” denotes the temporal-path
As shown in Figure 9, we give the detailed arches of the proposed Dual-path Spatial-Temporal Adapter (DSTA). For the spatial-path, we use the depth-wise spatial convolution with the kernel size of $1\times 1$ to capture the variation of micro-actions between consecutive frames. For the temporal-path, inspired by the bottleneck design in [52, 22], we design the depth-wise temporal convolution with the kernel size of $\{1,3,1\}$. After that, we use two separate parameters, $\beta$, and $\gamma$, to dynamically fuse the information from the spatial path and temporal path, respectively. The loss optimization of the proposed method is the same as the baseline model [41].
B.1 Evaluation Metric
Note that the evaluated methods can only predict the action-level categories. To make a fair comparison, we follow the common practice [17, 29, 16] directly mapping the predicted action-level category to body-level category. Then, we use the mAP [41] as the evaluation metric to evaluate the performance of the model on the action-level category and body-level category. To evaluate the performance at the action level and the body level simultaneously, we report the average value (AVG) derived from the mean of the body-level and action-level scores.
Appendix C More Experiments
Table 5: Ablation study of the position of the adapter.
       Layers     Action-level Detection-mAP
1-4 5-8 9-12 @0.2 @0.5 @0.7 Avg.
16.71 10.38 5.31 10.38
22.07 14.11 7.51 13.83
21.36 13.54 7.47 13.52
28.05 20.40 9.03 18.16
C.1 More Ablation Studies
Ablation studies on the position of the adapter. Here, we evaluate the effect of the position of the adapter. As shown in Table 5, we divided the 12 layers of VideoMAE-S into three stages, namely early (1-4), middle (5-8), and latter (9-12). The results show that the adapter on the middle and latter layers produces better results than the early layers. The adapter on all layers achieves the best result of 18.16 in average mAP.
(a) Action-level Results
(b) Body-level Results
Figure 10: Sensitivity Analysis. The left is the normalized mAP at tIoU=0.5, while the right is the relative normalized mAP change at tIoU=0.5.
Figure 11: Top-10 proposals of the proposed method. In the vertical axis, gt-$i$ represents the $i$-th ground truth, while pred-$i$ denotes the $i$-th micro-action proposal. The color bar denotes the tIoU between the proposal and ground truth actions.
C.2 Sensitive Analysis
As shown in Figure 10, we perform the sensitivity analysis of the proposed model. The Average-mAP results are categorized into different buckets by three characteristics, i.e., “Coverage”, “Length”, and “#Instances”. “Coverage” denotes the ratio of instances within the video, “Length” represents the duration (seconds) of instances, and “#Instances” is the number of instances. Following the False Negative analysis settings in the main paper, the range of these characteristics are as follows, i.e., “Coverage” refers to [0.0, 0.2, 0.4, 0.6, 0.8, 1.0], “Length” refers to [0, 3, 5, 8, 9, INF], and “#Instances” refers to [-1, 2, 5, 7, INF]. In each subplot, the left column represents the mAPN and average mAPN for each specific bucket on the MMAD dataset when tIoU=0.5, while the right graph summarizes the left graph by displaying the sensitivity range, calculated as the difference between the maximum and minimum mAPN values.
From the left column, we can find that: 1) In the characteristics of coverage and length, the average mAPN gradually increases. The lowest performance is the XS bucket and the highest is the L bucket. These results suggest that instances with more coverage or longer length are generally better recognized; 2) In the characteristics of “#Instances”, the average mAPN gradually decreases with the increase of action instances. The worst result is in the L bucket. These results indicate that dense instances will lead to lower performance. 3) The average mAPN at the body-level is better than the action-level, which suggests that the detection of action-level labels is easier. From the right column, we have drawn that “Coverage” and “Length” are the most sensitive factors, with a relative change of 1.7, while “#Instances” is the least sensitive factor, with a relative change of 1.1. These results indicate the number of instances has a small influence on performance.
In summary, future work should focus on improving the accuracy of recognizing the instances with shorter coverage and length, as these are the most sensitive factors. In addition, the Average-mAP scores at the body-level are generally higher than the action-level, implying that the body-level information can be used to enhance the detection accuracy at the action-level.
Appendix D More Visualization Results
As shown in Figure 11, we give more prediction results on the proposed MMA-52 dataset. In each subfigure, the red bars denote the ground truth instances, and the blue bars below them are the top 10 predicted proposals. The right color bar is the tIoU value between the proposal and the ground truth. The darker the color, the higher the value of tIoU. These results demonstrate that the proposed method achieves high precision in detecting micro-actions, even those with high co-occurrence and short durations.